OpenStudy (anonymous):

The first term of a sequence is 29, and the fourth term is 50. What are the second and third terms? 37 and 43 36 and 42 37 and 42 36 and 43

OpenStudy (anonymous):

i am going to guess that somewhere in the question it say "arithmetic sequence"

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

....

OpenStudy (anonymous):

something is wrong here

OpenStudy (anonymous):

are you sure this is the exact question no typos?

OpenStudy (anonymous):

oh nvm \(50-29=21\) and \(21\div 3=7\) add 7 each time

OpenStudy (anonymous):

\(29+7=36\) and \(36+7=43\) also \(43+7=50\)